write an equation in slope intercept form that passes through (-1,-5) and is parallel to the graph of y=-6x-4
Answer by nerdybill(7384)

write an equation in slope intercept form that passes through (-1,-5) and is parallel to the graph of y=-6x-4
.
slope of :
y=-6x-4
is -6
.
A parallel line will have the same slope (-6).
.
plug the slope (-6) and the point (-1,-5) into the "point-slope" form:
y - y1 = m(x - x1)
y - (-5) = -6(x - (-1))
y + 5 = -6(x + 1)
y + 5 = -6x - 6
y = -6x - 6 - 5
y = -6x - 11 (this is what they're looking for)